### v2

# A Language-Agnostic Hierarchical LoRA-MoE Architecture for CTC-based Multilingual ASR

Abstract
Large-scale multilingual ASR (mASR) models such as Whisper achieve strong performance but incur high computational and latency costs, limiting their deployment on resource-constrained edge devices. In this study, we propose a lightweight and language-agnostic multilingual ASR system based on a CTC architecture with domain adaptation. Specifically, we introduce a Language-agnostic Hierarchical LoRA-MoE (HLoRA) framework integrated into an mHuBERT-CTC model, enabling end-to-end decoding via LID-posterior-driven LoRA routing. The hierarchical design consists of a multilingual shared LoRA for learning language-invariant acoustic representations and language-specific LoRA experts for modeling language-dependent characteristics. The proposed routing mechanism removes the need for prior language identity information or explicit language labels during inference, achieving true language-agnostic decoding. Experiments on MSR-86K and the MLC-SLM 2025 Challenge datasets demonstrate that HLoRA achieves comparable performance to two-stage inference approaches while reducing RTF by 11.7% and 8.2%, respectively, leading to improved decoding efficiency for low-resource mASR applications.
publicationid: pubid: 0000–0000/00\$00.00 © 2021 IEEE
I Introduction
Large-scale multilingual speech recognition (mASR) models, such as Whisper [1] and SeamlessM4T [2], have achieved remarkable recognition performance by leveraging massive multilingual training data. However, the high computational cost and inference latency severely limit their practical deployment, especially on resource-constrained edge devices such as hearing aids and mobile platforms. This motivates interest in lightweight mASR systems that can balance recognition performance with efficiency and real-time requirements.
Most state-of-the-art mASR systems are built upon attention-based encoder–decoder (AED) architectures, with Whisper and its variants representing a dominant research direction [3]. These systems typically adopt byte-level byte-pair encoding (BBPE) [4] to unify token representations across languages and rely on large-scale multilingual data and extensive training resources. However, despite their impressive in-domain results, such models often suffer significant performance degradation under domain mismatch, including variations in speaking style, acoustic conditions, or accents that differ from the training data [5, 6, 7, 8].
To alleviate domain mismatch in multilingual ASR, recent studies have increasingly adopted parameter-efficient fine-tuning strategies, with low-rank adaptation (LoRA) [9, 10, 11]. Representative approaches [12, 13, 14, 15] adapt pretrained models to specific languages or speaking styles while largely preserving their original multilingual capabilities. Although these methods reduce adaptation costs, most existing approaches focus on monolingual [16, 17] or language-dependent LoRA fine-tuning [19, 20, 21, 18] and largely overlook the acoustic commonalities shared across languages. Moreover, many of these methods require explicit language identity (LID) during inference [22, 23, 24], which substantially restricts usability in real-world scenarios. In practical applications, users generally prefer language-agnostic mASR systems that do not require manual language selection between conversational turns.
Existing language-agnostic mASR approaches generally follow two paradigms. The first relies on universal modeling with shared vocabularies, such as unified BBPE token in Whisper [1]. While effective at large scale, such approaches often struggle to capture fine-grained language-specific acoustic characteristics. The second paradigm adopts two-stage inference pipelines, where an LID module first predicts the language and then activates a corresponding expert, such as router-based mixture-of-experts (MoE) systems [25, 26, 27]. Although effective, these pipelines introduce additional inference latency and are prone to error propagation from LID to ASR, making them less suitable for low-latency and on-device applications [28, 29, 30]. These limitations motivate a lightweight, end-to-end architecture for achieving language-agnostic behavior without sacrificing recognition performance or decoding efficiency.
Motivated by these challenges, this work focuses on CTC-based multilingual ASR and adopts mHuBERT-CTC as the backbone to develop a lightweight and truly language-agnostic framework with non-autoregressive, low-latency decoding. Our main contributions are summarized as follows:
  • A Language-Agnostic Hierarchical LoRA-MoE architecture is proposed and integrated into an mHuBERT-CTC framework, enabling multilingual ASR with single-pass, end-to-end decoding.
  • A hierarchical LoRA design combining a multilingual shared LoRA and language-specific LoRA modules is developed to jointly model language-invariant acoustic features and language-dependent characteristics, achieving improved efficiency and competitive accuracy.
  • An LID-posterior-driven LoRA routing mechanism is introduced, which dynamically activates language-specific LoRA experts from intermediate outputs and removes the need for prior LID information during inference.
  • A CTC-based language-agnostic baseline is built by adapting mHuBERT-CTC with language-specific LoRA modules and incorporating a two-stage inference strategy for systematic comparison.
Figure 1: The structure of LoRA-adapted mHubert-CTC mASR with two-stage inference (mHuBERT-CTC-LIDLoRA).
II LoRA-Adapted mHubert-CTC mASR with Two-Stage Inference
To establish a lightweight baseline for language-agnostic multilingual ASR, we first build a CTC-based system upon the mHuBERT-147 multilingual pretrained model [31]. With this backbone, language-specific LoRA modules are introduced for parameter-efficient adaptation, and a state-of-the-art two-stage inference strategy [13, 12] is adopted to enable language-agnostic decoding. The resulting system is referred to as mHuBERT-CTC-LIDLoRA. The overall architecture is illustrated in Fig. 1, and its components are described below.
mHuBERT-CTC backbone: To enable implicit language identification within a unified CTC framework, a dedicated language token is introduced as the first predict token in the CTC target sequence during training. This design allows the model to jointly learn LID and ASR in an E2E manner without an explicit LID module. The training procedure follows a standard fine-tuning strategy for SSL-pretrained models: first optimize the CTC layer while keeping the mHuBERT encoder frozen, and then fine-tune the entire network end-to-end.
LoRA-adapted mHuBERT-CTC: To improve multilingual adaptability, we adopt a LoRA-based fine-tuning strategy. Specifically, as shown in Fig. 1, independent LoRA modules are trained for each target language by freezing all pretrained mHuBERT-CTC parameters and updating only the language-specific LoRAs. This results in a multilingual ASR system with explicit language-dependent adaptation capability, while incurring a small number of additional trainable parameters.
Two-stage language agnostic inference: Language agnostic decoding is achieved through a two-stage inference procedure. In the first stage, the base mHuBERT-CTC model without LoRA adaptation is used to predict an LID token via CTC decoding, and the language with the highest posterior probability is selected. In the second stage, the corresponding language-specific LoRA module is activated to perform ASR decoding. Although this two-stage approach enables language-agnostic decoding without prior LID information, it introduces additional inference latency and is susceptible to error propagation from language prediction to ASR. These limitations motivate a more efficient single-pass and end-to-end solution.
III mHuBERT-CTC with Hierarchical LoRA-MoE Architecture
While language-specific LoRA adaptation enables parameter efficient multilingual ASR, independent per-language training is inefficient and fails to leverage cross-lingual acoustic commonalities. Existing two-stage language-agnostic inference further incurs redundant computation and increased latency, while relying on frozen LID predictions that cannot be improved during downstream adaptation. To address these issues, we propose a hierarchical LoRA-MoE architecture built upon mHuBERT-CTC (mHuBERT-CTC-HLoRA), which jointly optimizes LID and ASR, enables unified modeling of language-invariant and language-specific representations, and supports efficient single-pass, language-agnostic decoding.
Figure 2: The proposed mHuBERT-CTC-HLoRA architecture. The CNN and Transformer backbone are frozen, while the shared LoRA in the first $k$ Transformer layers, the language-specific LoRA in the last NkN\!-\!k layers and CTC head, as well as the LID linear layer, are trainable.
III-A Model Architecture
Fig. 2 illustrates the proposed mHuBERT-CTC-HLoRA architecture. An unknown-language speech is processed by a frozen CNN front-end and an $N$-layer Transformer encoder from the pretrained mHuBERT-147 model. To balance cross-lingual sharing and language-specific modeling, the encoder is divided into two parts: the lower $k$ layers incorporate a shared LoRA optimized across all languages to capture language-invariant acoustic representations, while the upper NkN-k layers employ language-specific LoRA modules to model discriminative language-dependent characteristics. Intermediate representation after $k$-th layer is fed into a LID classifier, whose posterior dynamically activates the corresponding language-specific LoRA in upper layers and CTC head, thereby tightly coupling LID and ASR within a unified framework and enabling single-pass, language-agnostic decoding.
III-B Model Training
Let the multilingual training set be denoted as $D=\{D_{1},D_{2},\dots,D_{L}\}$, where $D_{i}$ contains utterances from the $i$-th language. Each language-specific subset is first split into minibatches. All minibatches across languages are then pooled together and randomly shuffled at the beginning of every training epoch. At each iteration, a minibatch from language $i$ is randomly sampled. The input speech $X$ is processed by the frozen CNN front-end and the first $k$ Transformer layers augmented with the shared LoRA:
$X_{h}=\mathcal{F}_{1}\left(\texttt{CNN}\left(X\right),W_{s},\theta_{trans}^{k}\right)$ (1)
where $W_{s}$ denotes the shared LoRA parameters and $\theta^{k}_{\text{trans}}$ are the frozen pretrained weights. Next, $X_{h}$ is forwarded to the remaining NkN-k Transformer layers with the language-specific LoRA $W_{L}^{i}$, selected according to the ground-truth language label:
$X_{N}=\mathcal{F}_{2}\left(X_{h},W_{L}^{i},\theta_{trans}^{N-k}\right)$ (2)
followed by CTC decoding to compute the ASR loss:
$\mathcal{L}_{ASR}=\texttt{CTC}\left(X_{N},W_{L}^{i}\right)$ (3)
In parallel, $X_{h}$ is fed into the LID classifier(a projection layer) to predict the language:
$Y_{L}$ $=\texttt{LIDLinear}\left(X_{h}\right)$ (4)
$\mathcal{L}_{LID}$ $=\texttt{CE}\left(Y_{L},Y_{L}^{true}\right)$ (5)
The total training objective is defined as a weighted combination of ASR and LID losses with $\lambda$:
$\mathcal{L}=(1-\lambda)\mathcal{L}_{ASR}+\lambda\mathcal{L}_{LID}$ (6)
Although only one language-specific LoRA $W_{L}^{i}$ is activated per iteration, the shared LoRA $W_{s}$ and the LID classifier are updated in every iteration. By alternating languages through shuffled multilingual minibatches, the proposed training strategy enables balanced optimization of all language-specific LoRA modules while jointly learning shared acoustic representations and language discrimination.
III-C Inference strategy
During inference, the proposed mHuBERT-CTC-HLoRA framework performs single-pass language-agnostic decoding via an LID-posterior-driven LoRA routing mechanism. Given an input utterance, the shared LoRA modules in the lower Transformer layers are first applied to extract language-invariant representations. Based on the intermediate outputs, an integrated LID classifier predicts a posterior distribution over the target languages. This LID posterior is then directly used for hard argmax routing to activate the corresponding language-specific LoRA expert in the upper Transformer layers and the CTC head within the same forward pass, producing the final transcription.
Unlike the two-stage inference baseline described in Section II, which relies on a separate encoder pass and external LID decision, the proposed approach jointly performs language identification and ASR in a unified, end-to-end manner. By leveraging LID posteriors estimated from intermediate shared representations, HLoRA eliminates the need for prior LID information, alleviates error propagation across stages, and significantly reduces inference latency. As a result, the proposed inference strategy enables efficient, low-latency, and truly language-agnostic multilingual ASR.
IV Experiments and Results
IV-A Datasets
Experiments are conducted on the MSR-86K [32] and the MLC-SLM 2025 Challenge [5] datasets. MSR-86K is used to build the mHuBERT-CTC pretrained ASR model and taken as source-domain mASR task. From its 1500-hour subset, 11 languages are selected: English (EN), French (FR), German (DE), Italian (IT), Japanese (JA), Korean (KO), Portuguese (PT), Russian (RU), Spanish (ES), Thai (TH), and Vietnamese (VI). MLC-SLM dataset is taken as the low-resource target domain mASR dataset for system evaluation, focusing on Indian-accented English (EN-IN), FR, DE, JA, and KO. For each language, 100 hours of audio are used for training (98h) and development (2h), while the official development and test sets provided by the challenge are used for evaluation.
IV-B Configurations
We adopt the pretrained SSL model mHuBERT-147 [31] as the backbone. SentencePiece tokenization uses character units for JA, KO, and TH and a 4k BPE vocabulary for other languages, resulting in a unified vocabulary of 9,521 tokens. LoRA is applied to the self-attention Q, K, V projections and the CTC layer with rank $r=32$ and $\alpha=64$. Evaluation is performed using mixture error rate, with character error rate (CER) for JA/KO/TH and WER otherwise).
TABLE I: Overall mASR results on MLC-SLM datasets. MSR-1500h and MLC-500h denote training data subsets selected from MSR-86K and MLC-SLM, respectively. $\surd$” and $\times$” indicate language-known and language-agnostic inference, while “single” and “double” denote single-pass and two-stage inference, respectively. “RTF” indicates real-time factor measured on a single NVIDIA RTX 4090 GPU(24G) with beam size 10 and batch size 1. For S1–S6, all systems share the same mHuBERT-147 backbone pretrained on 90k hours of multilingual data.
ID System Params(M) TrainingData LID Inference RTF Evaluation Sets(WER%)
MLC-dev MLC-test MLC-dev MLC-test
S1 mHuBERT-CTC(baseline) 97 MSR-1500h × single 0.553 0.515 44.5 43.7
S2 MLC-500h × single 0.549 0.514 22.5 21.6
S3 mHuBERT-CTC-LIDLoRA 107 MLC-500h $\surd$ single 0.563 0.539 24.6 23.0
S4 × double 0.592 0.564 26.6 24.8
S5 mHuBERT-CTC-HLoRA 102 MLC-500h $\surd$ single 0.526 0.531 26.0 24.0
S6 × single 0.523 0.518 26.3 24.7
S7 Whisper-base [1] 74 680,000h × single - - 37.2 35.4
S8 Whisper-small [1] 244 680,000h × single - - 29.7 28.6
IV-C Overall Results
Table I reports overall mASR performance on the MLC-SLM 2025 development and test sets. S1 is trained on source-domain MSR-1500h, while S2 initializes from S1 and is fully fine-tuned on target-domain MLC-500h. The large gap between S1 and S2 highlights the severe impact of domain mismatch on conversational speech. Moreover, state-of-the-art Whisper models (S7 and S8) still underperform S2 on the test set, emphasizing the need for efficient domain adaptation in pretrained mASR systems.
S3 and S4 evaluate LoRA-based parameter-efficient adaptation on the source-domain mHuBERT-CTC baseline (S1), updating only a small subset of parameters. As expected, this parameter-efficient setting leads to a noticeable performance gap compared with full fine-tuning in S2. However, this behavior is reasonable in low-resource adaptation scenarios, where full fine-tuning of large pretrained models is computationally expensive and prone to overfitting to the target domain, potentially degrading performance on the source domain.
S3 uses language-dependent inference, activating the language-specific LoRA module with the ground-truth label, while S4 performs language-agnostic decoding via a two-stage pipeline. Although S4 achieves fully language-agnostic operation, its recognition performance is consistently lower than that of S3. This gap primarily arises from the two-stage design, where the activation of language-specific LoRA modules relies on the LID prediction produced by the base mHuBERT-CTC model. Since this LID component is inherited from a frozen pretrained model and cannot be jointly optimized during LoRA adaptation, errors in language prediction directly affect subsequent ASR decoding. These observations reveal a fundamental limitation of two-stage language-agnostic inference and motivate the proposed single-pass HLoRA framework, where LID prediction and ASR are jointly optimized end-to-end.
S5 and S6 evaluate HLoRA under language-known and language-agnostic inference, respectively. S5 shows slightly degraded performance than S3 due to reduced language-specific LoRA capacity. By design, HLoRA allocates fewer language-specific parameters in order to enable a lightweight and streaming-friendly architecture, trading a small amount of accuracy for improved efficiency and unified modeling.
More importantly, S6 demonstrates superior efficiency by achieving competitive performance via unified single-pass inference. S6 achieves competitive WERs, which are comparable to or slightly better than the two-stage baseline S4, while reducing relative RTF by 11.7% and 8.2%. The critical advantage lies in the HLoRA framework’s ability to jointly optimize LID and ASR within a single, unified model, thereby enabling single-pass inference. This design not only mitigates the computational overhead and latency associated with two-stage inference but also confirms that robust language-aware adaptation can be achieved implicitly, without relying on explicit LID input at inference time. These results validate HLoRA as an efficient and practical solution for low-latency, language-agnostic multilingual ASR.
IV-D Ablation Results
TABLE II: mHuBERT-CTC-HLoRA ablation study with different $k$. Language-wise results are on the MLC-dev/MLC-test in WER%.
Language $k=1$ 3 6 9 11
English 40.5/48.1 33.8/36.6 29.0/26.6 29.2/25.2 28.9/25.3
French 26.6/33.9 26.3/33.2 26.1/31.7 26.4/32.0 26.3/32.2
German 61.2/54.3 54.9/44.1 39.1/32.9 39.5/32.6 39.8/33.4
Japanese 37.3/41.0 32.6/37.0 25.0/27.8 25.5/23.6 25.7/24.7
Korean 17.9/16.5 17.5/16.0 17.2/16.1 17.7/16.6 18.1/16.9
avg 34.8/36.6 31.3/32.1 26.0/26.0 26.3/24.7 26.5/25.2
Table II presents the ablation results of the proposed mHuBERT-CTC-HLoRA model with different numbers of shared Transformer layers $k$. The results reveal a clear trade-off between cross-lingual sharing and language-specific adaptation. When $k$ is small (e.g., $k=1$ or $k=3$), most Transformer layers rely on language-specific LoRA modules, leading to insufficient cross-lingual sharing and inferior performance across all languages. As $k$ increases, the model benefits from stronger language-invariant representations learned by the shared LoRA, resulting in substantial WER reductions. The best overall performance is achieved with a moderate number of shared layers ($k=6$ or $k=9$), which yields the lowest average WER on both MLC-dev and MLC-test. Further increasing $k$ to 11 slightly degrades performance, suggesting that excessive parameter sharing limits the model’s ability to capture language-specific acoustic characteristics. These results confirm that an appropriate hierarchical decomposition of shared and language-specific LoRA modules is critical for balancing model capacity, and efficiency.
In addition, Fig. 3 further compares the LID behavior of LIDLoRA and the proposed HLoRA. HLoRA produces more diagonal-dominant confusion matrices across all five target languages, leading to an average LID accuracy improvement from 90.1% to 97.9%. This gain is attributed to the single-pass, end-to-end HLoRA design, which jointly optimizes LID and ASR using shared LoRA features, avoiding the frozen and non-adaptive LID stage in two-stage inference and enabling more reliable language-agnostic decoding.
Figure 3: Confusion matrix of LID results on the MLC dev-set for LIDLoRA (left) and the proposed HLoRA (right). “OTS” denotes the aggregated predictions of the remaining six source-domain languages in the pretrained mHuBERT-CTC model that are not included in five target-domain languages.
V Conclusion
This paper presented HLoRA, a lightweight and streaming-friendly multilingual ASR framework designed for low-resource, language-agnostic scenarios. By hierarchically decomposing LoRA modules into shared and language-specific components, the proposed method effectively balances cross-lingual representation sharing and language-dependent acoustic modeling. Extensive experiments on the MLC-SLM 2025 benchmark demonstrate that HLoRA consistently outperforms strong two-stage language-agnostic baselines while avoiding external LID supervision. Moreover, the single-pass design significantly improves LID robustness and inference efficiency, making the system more suitable for practical multilingual deployment. Our models and source code are publicly released at https://github.com/zhengyuang7/HLoRA to facilitate reproducibility and future research.